Evaluate the function for the value specified – you need to arrive at a final numerical answer for each part. f(x) = x2 – 7x + 10
Find:
a) f (0)
b) f (5

c) f (-2)

Answer by stanbon(57361)

f(x) = x^2 – 7x + 10
Find:
a) f (0)= 0^2-7(0)+10 = 10
b) f (5)= 5^2-7(5)+10 = 0
c) f(-2)= (-2)^2 -7(-2)+10= 28
Cheers,
Stan H.